OpenStudy (anonymous):

What is 5/8 of 32/115 of 161/200?

Parth (parthkohli):

of means to multiply. \(\Large \color{Black}{\Rightarrow {5 \over 8} \times {32 \over 115} \times {161 \over 200} }\) Now a calculator.

OpenStudy (lgbasallote):

\[\large \frac{5}{8} \times \frac{32}{115} \times \frac{161}{200}\]

Parth (parthkohli):

My calculator gets 0.14 or \(14 \over 100\)

OpenStudy (lgbasallote):

are you sure o.O

Parth (parthkohli):

Cross my heart.

Parth (parthkohli):

Or alternatively you can also cancel.

OpenStudy (anonymous):

not 7/50?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {7 \over 50} = {14 \over 100} = {0.14}}\) All are equivalent.

Parth (parthkohli):

So you were right.

OpenStudy (lgbasallote):

7/50 is right :D

OpenStudy (anonymous):

Thanks :) hahaha.